A motorboat takes 4 hours to travel 128km going upstream. The return trip takes 2 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?
Answer by macston(5194)

r=rate of boat in still water, c=current
128 km/ 4hours=r-c=32 km/hr
128 km/2 hours=r+c=64 km/hr
r-c=32 km/hr add c to each side
r=32 km/hr+c
r+c=64 km hr Substitute for r
32 km/hr + c + c=64 km/hr Subtract 32 km/hr from each side.
2c=32 km/hr Divide each side by 2.
c=16 km/hr ANSWER 1: The current is 16 km/hr
r=32 km/hr + c=32 km/hr+ 16 km/hr=48 km/hr
ANSWER 2: The speed of the boat in still water would be 48 km/hr.